# Neural Networks and Friction: Slide, Hold, Learn

### 21 Feb 2024 v1

@@ Abstract
In this study, it is demonstrated that Recurrent Neural Networks (RNNs), specifically those utilizing Gated Recurrent Unit (GRU) architecture, possess the capability to learn the complex dynamics of rate-and-state friction laws from synthetic data. The data employed for training the network is generated through the application of traditional rate-and-state friction equations coupled with the aging law for state evolution. A novel aspect of our approach is the formulation of a loss function that explicitly accounts for initial conditions, the direct effect, and the evolution of state variables during training. It is found that the RNN, with its GRU architecture, effectively learns to predict changes in the friction coefficient resulting from velocity jumps, thereby showcasing the potential of machine learning models in understanding and simulating the physics of frictional processes. The implications of these findings for the field of geophysics and fault mechanics are discussed, highlighting the model’s proficiency in capturing the intricate behavior of seismic slip phenomena.
@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [4].
Friction modeling [5, 6] aims to capture essential features observed in experimental settings [7], notably the direct effect and healing [2]. The direct effect, also known as the velocity strengthening behavior, refers to the immediate increase in frictional resistance with an increase in sliding velocity. This phenomenon can be described by the empirical equation [4]:
@@
where $D_{c}$ is a critical slip distance over which the state variable evolves significantly [9]. The state is interpreted as a measure of effective age of contacts between the asperities of the opposed surfaces [8], and it therefore has units of time.
@@
Rate-and-state friction (RSF) laws stand out by explicitly addressing the direct effect and healing, thereby providing a more comprehensive framework for understanding the seismic cycle [11]. These laws are formulated as:
@@
Criticism has been levied against the use of the ”state” variable in RSF laws for its empirical nature and lack of a clear physical basis, leading to calls for models that include more explicit physical mechanisms [13, 14]. To begin, the parameters that appear in eq. 3 are assumed to be constants, and they may not [15, 16, 17, 18]. Furthermore, the need to regularize these laws to address numerical instabilities and ensure the physical realism of simulations has been highlighted [19, 20]. Recently, laboratory [21] and virtual experiments [22, 23] have shown the limitations of RSF laws even in highly controlled interface conditions.
@@
This success in modeling path-dependent phenomena in materials science provides a strong foundation for the belief that similar methodologies can be applied with success to the study of friction. Friction, much like material plasticity, is inherently dependent on the history of contact and slip between surfaces, characterized by complex state-dependent behaviors such as the direct effect and healing. The capability of RNNs to capture and learn from sequential, history-dependent data makes them promising candidates for advancing our understanding and modeling of friction dynamics. Their ability to learn from data where previous states influence future behavior aligns well with the challenges presented by rate-and-state friction laws, suggesting that these neural network models could offer new insights and predictive capabilities beyond traditional friction models.
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular. This approach is essentially different from the one used recently by Ishiyama and collaborators [37], who used machine learning (random forest [38]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[14, 39, 40]. It also differs from the use made by Wang et al [41] and Karimpouli et al [42], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without internal variables, since the state is not a physical quantity that can be measured experimentally (assuming it even exists as such). In reality, the complexities of interface physics [43] are such that probably no simplified internal variable can be defined as long as any real one aims at capturing a realistic degree of complexity We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
@@ II.1 Neural networks for sequences: gated recurrent units
RNNs are engineered to handle sequential data thanks to their ability to maintain a memory of previous inputs by feeding the output of a neuron back onto itself [25]. This recursive nature allows making predictions based on the information accumulated from prior inputs. However, standard RNNs often struggle with long-term dependencies due to the vanishing gradient problem, where gradients become too small or too large for the network to learn effectively over many time steps.
Gated Recurrent Units (GRUs) are an evolution of the traditional RNN architecture, designed to mitigate the vanishing gradient problem and improve the network’s ability to learn from long sequences. GRUs introduce gating units that control the flow of information. These gates “decide” which information is relevant to keep or discard as the sequence progresses, enabling the network to maintain a longer memory and capture dependencies over longer sequences.
@@
    $\tilde{h}_{t}=\tanh(Wx_{t}+U(r_{t}\odot h{t-1})+b)\,.$ (6)
@@
In these equations, $x_{t}$ represents the input at time step $t$; $h_{t}$ is the hidden state at time $t$; $\sigma$ denotes the sigmoid activation function (to ensure the gate activations are constrained between 0 and 1); $\tanh$ is the hyperbolic tangent function (introducing non-linearity); $W$, $U$, and $b$ are the weights and biases associated with the input and hidden states, respectively; and $\odot$ signifies element-wise multiplication.
@@ II.2 Synthetic data generation
As mentioned earlier, we do not consider the state $\theta$ to be a necessary variable, thus it will not be used henceforth.
To generate the dataset for training and verifying our model, we start by defining a traditional RSF model with specific parameter values, grounded in the fundamental equations presented earlier (eq. 3), with parameter values set as follows:
  • Friction coefficient, $\mu_{0}=0.5$
@@
  • Reference velocity, $V_{\text{ref}}=1.0\times 10^{-4}$ m/s
  • Characteristic slip distance, $D_{c}=0.2\times 10^{-3}$ m
Using these values, we simulate 15000 random sequences to reflect diverse frictional sliding scenarios. Each sequence is designed to simulate a total sliding distance of $30D_{c}$ and is completed within 0.013 seconds. The number of velocity jumps within each sequence is randomly determined, drawing from a uniform distribution with values 3, 4, 5. It is assumed that each sequence initiates from a non-zero velocity, with the initial state of the interface corresponding to the steady state for that velocity, as defined by the RSF laws: $\theta_{ss}=D_{c}/V_{0}$.
Velocity jumps within these sequences occur at random intervals, determined by sampling times from a uniform distribution across the interval 0, 1, with 0 marking the start and 1 the end of the sliding protocol. Each sequence incorporates a hold period where the velocity drops to zero, simulating intervals where sliding stops until the subsequent jump.
To compute the velocities between jumps, we divide the total sliding distance by the time spans between these jumps, ensuring that the velocities remain constant in these intervals. Thus, we define the “velocity protocol” for each sequence, i.e., virtual experiment.
Having the velocity protocol, we proceed to calculate the evolution of the state variable $\theta$ using eq. 2 and initial condtions $V(t=0)=V_{0}$ and $\theta(t=0)=\theta_{ss}$. With the velocity protocol and the evolving state variable at hand, we can then compute the friction coefficient’s evolution over time, as described by the RSF model (eq. 3).
The dataset comprises pairs of “features” and “targets”. Features represent the velocity protocol sampled at 100 equally-spaced instants during the duration of each virtual experiment, while targets correspond to variations of the friction coefficient sequence at these same instants. What is the reason for working with variations and not with the actual values of friction coefficient? There is no way to predict the initial values of friction based on these datasets, they have to be provided as initial conditions over which the variations are superimposed.
@@ II.3 Network details
Mathematica’s neural networks package has been used.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation. Despite its simplicity, this network combines memory and transformation capabilities, make it well-suited to our goal of understanding and predicting the dynamic behavior of rate-and-state friction laws from sequential data.
@@
II.3.2 Training hyperparameters
We employ the ADAM optimization method due to its adaptive learning rate capabilities, which helps in converging to the optimal solution more quickly and efficiently than conventional gradient descent methods. A learning rate of 0.001 is chosen as a starting point because it strikes a balance between fast convergence and the risk of overshooting the minimum of the loss function. This value is generally considered a good default and has been proven effective across a wide range of tasks. In order to ensure that each training batch has enough data to represent the dataset’s diversity while keeping the computational demand manageable, we choose to use a batch size of 32. This size is large enough to benefit from computational optimizations during training yet small enough to avoid the potential pitfalls of training with very large batches, such as poor generalization. Finally, we use 50 epochs to allow the model opportunity to learn from the data without excessive training time. This limit helps prevent overfitting by avoiding too many epochs over the training dataset, promoting a model that generalizes well to unseen data.
@@ II.3.3 Loss function
Inspired by physics-informed neural networks, we devise a loss function to guide the optimization of the parameters of the network. The conditions that observations impose over the friction coefficient are the direct effect, i.e., logarithmic dependence of velocity magnitude, and logarithmic healing, i.e., logarithmic increase of the static friction coefficient with time (during holds)
@@
Let $\delta\bm{\mu}$ and $\delta\bm{\hat{\mu}}$ to represent the vector of normalized predictions and targets (respectively) for a given vector of features (velocity protocol) $\bm{V}$, i.e., $\delta\bm{\mu}=\mathsf{NN}(\bm{V}/V_{ch})$. See that $\delta\bm{\mu},\,\delta\bm{\hat{\mu}},\,\bm{V}\in\mathbb{R}^{N}$ (and $N=100$ in this case). The first part of the loss is the mean-squared error:
@@
where |||||$\cdot$||| | ⋅ | | represents the $L_{2}$ norm of the vector and $N=100$ is the number of elements in the sequence. In overall graph for training, fig. 6, this operation is implemented by the neuron labeled $\mathsf{MS}$.
@@
where the sub-index $1$ refers to the first element of every sequence. The implementation of this operation is depicted in fig. 3: it takes the first sequence value of both targets (Targets) and predictions (Predictions), subtract each other and square the difference. Thus, the output of the net is eq. 10.
@@
The third term of the loss is meant to enforce the direct effect: a sudden peak in friction induced by a jump in leading velocity (which, moreover, is known to be proportional to the logarithm of the ratio of velocities before and after the jump, see section and I and appendix A). Assume that there are $N_{J}$ jumps in a given protocol, the term that measures the error between the one predicted by the network and the real one is
@@
the symbol $i$ indexes each jump, while $J_{i}$ refers to the entry of the vector where the jump begins, hence the meaning of the notation $\Delta(\delta\mu)_{J_{i}}=\delta\mu_{J_{i}+1}-\delta\mu_{J_{i}}$. The network depicted in fig. 4 implements this procedure: the inputs are the predictions generated by the GRU net (fig. 2) Predictions, the indices of the entries where velocoity jumps begin PosBeforeJ and after PosAfterJ, and the real magnitude of the friction $\mathsf{\Delta F_{data}}$, the output is the evaluation of eq. 11.
@@
Finally, interface strengthening during hold periods (a.k.a. healing) is recognized to happen logarithmically with respect to hold time, i.e., counting from the moment that the interface ceases to move. Consider that, given a sequence, the hold period starts at the $j$-th entry, and it lasts until $(j+N_{h})$-th entry. We can gather these terms in a separate vector $\delta\bm{\mu}_{h}=\left\{\delta\mu_{n}\right\}_{i=k}^{i=k+N_{h}}=\left\{\delta\mu_{n}\right\}_{n=1}^{n=N_{h}}$, where the sub-index $n$ has been introduced to index the values during hold straightforwardly. So the healing condition can be expressed as:
@@
Thus, the loss terms is:
@@
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. We found $\alpha=1$, $\beta=0.1$ and $\gamma=5\cdot 10^{-3}$ to yield satisfactory results, yet probably not optimal (see discussion section).

### 23 Feb 2024 v2

@@ Abstract
In this study, it is demonstrated that Recurrent Neural Networks (RNNs), specifically those utilizing Gated Recurrent Unit (GRU) architecture, possess the capability to learn the complex dynamics of rate-and-state friction laws from synthetic data. The data employed for training the network is generated through the application of traditional rate-and-state friction equations coupled with the aging law for state evolution. A novel aspect of our approach is the formulation of a loss function that explicitly accounts for initial conditions, the direct effect, and the evolution of state variables during training. It is found that the RNN, with its GRU architecture, effectively learns to predict changes in the friction coefficient resulting from velocity jumps, thereby showcasing the potential of machine learning models in understanding and simulating the physics of frictional processes.
@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3, 4]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [5].
Friction modeling [6, 7] aims to capture essential features observed in experimental settings [8], notably the direct effect and healing [2]. The direct effect refers to the immediate change in frictional resistance to a sudden change in sliding velocity. This phenomenon can be described by the empirical equation [5]:
@@
where $D_{c}$ is a critical slip distance over which the state variable evolves significantly [10]. The state is interpreted as a measure of effective age of contacts between the asperities of the opposed surfaces [9, 11], and it therefore has units of time. Equation 2 is termed “aging law”.
@@
Rate-and-state friction (RSF) laws stand out by explicitly addressing the direct effect and healing, thereby providing a more comprehensive framework for understanding the seismic cycle [13]. A customary expression of this law is [5]:
@@
Criticism has been levied against the use of the “state” variable in RSF laws for its empirical nature and lack of a clear physical basis, leading to calls for models that include more explicit physical mechanisms [15, 16]. To begin, the parameters that appear in eq. 3 are assumed to be constants, and they may not [17, 18, 19, 20, 21]. Furthermore, the need to regularize these laws to address numerical instabilities and ensure the physical realism of simulations has been highlighted [22, 23]. Recently, laboratory [24] and virtual experiments [25, 26] have shown the limitations of RSF laws even in highly controlled interface conditions.
@@
This success in modeling path-dependent phenomena in materials science provides a strong foundation for the belief that similar methodologies can be applied with success to the study of friction. Friction, much like material plasticity, is inherently dependent on history, in its case of contact and slip between surfaces. The capability of RNNs to capture and learn from sequential, history-dependent data makes them promising candidates for advancing our understanding and modeling of friction dynamics.
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular, eq. 2. This approach is essentially different from the one used recently by Ishiyama and collaborators [40], who used machine learning (random forest [41]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[16, 42, 43]. It also differs from the use made by Rouet-Leduc et al [44] and Hulbert et al [45], among others [46, 47], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without internal variables, since the state is not a physical quantity that can be measured experimentally (assuming it even exists as such). In reality, the complexities of interface physics [48] are such that probably no simplified internal variable can be defined as long as one aims at capturing any modestly ambitious degree of complexity. We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated by the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
@@ II.1 Neural networks for sequences: gated recurrent units
RNNs are engineered to handle sequential data thanks to their ability to maintain a memory of previous inputs by feeding the output of a neuron back onto itself [28]. This recursive nature allows making predictions based on the information accumulated from prior inputs. However, standard RNNs often struggle with long-term dependencies due to the vanishing gradient problem [30], where gradients become too small or too large for the network to learn effectively over many time steps.
Gated Recurrent Units (GRUs) are an evolution of the traditional RNN architecture, designed to mitigate the vanishing gradient problem and improve the network’s ability to learn from long sequences. GRUs introduce gating units that control the flow of information. These gates “decide” which information is relevant to keep or discard as the sequence progresses, enabling the network to maintain a longer memory.
@@
    $\tilde{h}_{t}=\tanh(Wx_{t}+U(r_{t}\odot h{t-1})+\xi)\,.$ (6)
@@
In these equations, $x_{t}$ represents the input at time step $t$; $h_{t}$ is the hidden state at time $t$; $\sigma$ denotes the sigmoid activation function (to ensure the gate activations are constrained between 0 and 1); $\tanh$ is the hyperbolic tangent function (introducing non-linearity); $W$, $U$, and $\xi$ are the weights and biases associated with the input and hidden states, respectively; and $\odot$ signifies element-wise multiplication.
@@ II.2 Synthetic data generation
As mentioned earlier, it is assumed that the state $\theta$ will not be needed by the neural network that is to replace the traditional RSF law. This means that there will not be vectors of values of $\theta$ being used during training and that the network will not , but it does not imply that eq. 2 is not being used during the datasets’ generation. Let us reiterate that we assume a hypothetical interface that is perfectly described by eq. 3 in tandem with eq. 7, but this fact would be hidden to us if we were to analyze experimental data coming out of that interface.
So, to generate the dataset for training and verifying our model, we start by defining a traditional RSF model with specific parameter values, grounded in the fundamental equations presented earlier (eq. 3), with parameter values set as follows:
  • Friction coefficient, $\mu_{0}=0.5$ (unitless)
@@
  • Reference velocity, $V_{\text{ref}}=1.0\cdot 10^{-4}$ m/s
  • Characteristic slip distance, $D_{c}=2.0\cdot 10^{-4}$ m
Using these values, we simulate 15000 random sequences to reflect diverse frictional sliding scenarios. Each sequence is designed to simulate a total sliding distance of $30D_{c}$ and is completed within 0.013 seconds. The number of velocity jumps within each sequence is randomly determined, drawing from a uniform distribution with values 2, 3, 4, 5. It is assumed that each sequence initiates from a non-zero velocity, with the initial state of the interface corresponding to the steady state for that velocity, as defined by the RSF laws: $\theta_{ss}=D_{c}/V(t=0)$.
Velocity jumps within these sequences occur at random intervals, determined by sampling times from a uniform distribution across the interval 0, 1, with 0 marking the start and 1 the end of the sliding protocol. Each sequence incorporates a hold period where the velocity drops to zero, simulating intervals where sliding stops until the subsequent jump. To avoid errors associated to dividing by zero or trying to compute the logarithm of zero, a tiny yet finite velocity is assigned to the hold for computations ($2\cdot 10^{-9}$ m/s).
To compute the velocities between jumps, we divide the total sliding distance by the time spans between these jumps, ensuring that the velocities remain constant in these intervals. Thus, we define the “velocity protocol” for each sequence / virtual experiment.
Having the velocity protocol, we proceed to calculate the evolution of the state variable $\theta$ using eq. 2 and initial conditions $V(t=0)=V_{0}$ and $\theta(t=0)=\theta_{ss}$. With the velocity protocol and the evolving state variable at hand, we can then compute the friction coefficient’s evolution over time, as described by the RSF model (eq. 3).
The dataset comprises pairs of “features” and “targets”. Features represent the velocity protocol sampled at 100 equally-spaced instants during the duration of each virtual experiment, while targets correspond to variations of the friction coefficient sequence at these same instants. What is the reason for working with variations and not with the actual values of friction coefficient? There is no way to predict the initial values of friction based on these datasets, they have to be provided as initial conditions over which the variations are later superimposed.
@@ II.3 Network details
Mathematica’s [49] neural networks package has been used.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation.
@@
II.3.2 Training parameters
We employ the ADAM optimization method [50] due to its adaptive learning rate capabilities, which helps in converging to the optimal solution more quickly and efficiently than conventional gradient descent methods. A learning rate of 0.001 is chosen as a starting point because it strikes a balance between fast convergence and the risk of overshooting the minimum of the loss function. This value is generally considered a good default and has been proven effective across a wide range of tasks. In order to ensure that each training batch has enough data to represent the dataset’s diversity while keeping the computational demand manageable, we choose to use a batch size of 32. Finally, we use 50 epochs to allow the model opportunity to learn from the data without excessive training time.
@@ II.3.3 Loss function
Inspired by physics-informed neural networks (PINNs) [32], we devise a loss function that reflects physical principles to guide the optimization of the parameters of the network. The conditions that observations impose over the friction coefficient are the direct effect, i.e., logarithmic dependence of velocity magnitude, and healing, i.e., logarithmic increase of the static friction coefficient with time (during holds)
@@
Let $\delta\bm{\mu}$ and $\delta\bm{\hat{\mu}}$ represent the vector of normalized predictions and targets (respectively) for a given vector of features (velocity protocol) $\bm{V}$, i.e., $\delta\bm{\mu}=\mathsf{NN}(\bm{V}/V_{ch})$. See that $\delta\bm{\mu},\,\delta\bm{\hat{\mu}},\,\bm{V}\in\mathbb{R}^{N}$ (and $N=100$ in this case). The first part of the loss is the mean-squared error:
@@
where |||||$\cdot$||| | ⋅ | | represents the $L_{2}$ norm of the vector. In overall graph for training, fig. 6, this operation is implemented by the neuron labeled $\mathsf{MS}$.
@@
where the sub-index $1$ refers to the first element of every sequence. The implementation of this operation is depicted in fig. 3: it takes the first sequence value of both targets (Targets) and predictions (Predictions), subtracts each other and squares the difference. Thus, the output of the net is eq. 10.
@@
The third term of the loss is meant to enforce the direct effect: a sudden peak in friction induced by a jump in leading velocity (which, moreover, is known to be proportional to the logarithm of the ratio of velocities before and after the jump, see section I and appendix A). Assume that there are $N_{J}$ jumps in a given protocol, the term that measures the error between the one predicted by the network and the real one is
@@
the symbol $i$ indexes each jump, while $J_{i}$ refers to the entry of the vector where the jump begins, hence the meaning of the notation $\Delta(\delta\mu)_{J_{i}}=\delta\mu_{J_{i}+1}-\delta\mu_{J_{i}}$. The network depicted in fig. 4 implements this procedure: the inputs are the predictions generated by the GRU net (fig. 2) Predictions, the indices of the entries where velocity jumps begin PosBeforeJ and after PosAfterJ, and the real magnitude of the friction jump $\mathsf{\Delta F_{data}}$, the output is the evaluation of eq. 11.
@@
Finally, interface strengthening during hold periods (a.k.a. healing) is recognized to happen logarithmically with respect to hold time, i.e., counting from the moment that the interface ceases to move. Consider that, given a sequence, the hold period starts at the $j$-th entry, and it lasts until $(j+N_{h})$-th entry. We can gather these terms in a separate vector $\delta\bm{\mu}_{h}=\left\{\delta\mu_{i}\right\}_{i=k}^{i=k+N_{h}}=\left\{\delta\mu_{n}\right\}_{n=1}^{n=N_{h}}$, where the sub-index $n$ has been introduced to index the values during hold straightforwardly. So the healing condition can be expressed as:
@@
Thus, the loss term is:
@@
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. It has been found that $\alpha=1$, $\beta=0.1$ and $\gamma=5\cdot 10^{-3}$ yields satisfactory results, yet probably not optimal (see discussion section).